Internet Engineering Task Force

INTERNET-DRAFT  Schulzrinne
draft-schulzrinne-sip-urls-00.ps  Columbia U.

April 22, 2001

Expires: December 2001

SIP URLs for Common Services, Roles and Functions

Status of this Memo

This document is an Internet-Draft and is in full conformance with all provisions of Section 10 of RFC2026.

Internet-Drafts are working documents of the Internet Engineering Task Force (IETF), its areas, and its working groups. Note that other groups may also distribute working documents as Internet-Drafts.

Internet-Drafts are draft documents valid for a maximum of six months and may be updated, replaced, or obsoleted by other documents at any time. It is inappropriate to use Internet-Drafts as reference material or to cite them other than as “work in progress.”

The list of current Internet-Drafts can be accessed at http://www.ietf.org/ietf/1id-abstracts.txt To view the list Internet-Draft Shadow Directories, see http://www.ietf.org/shadow.html.

Copyright Notice

Copyright (c) The Internet Society (2001). All Rights Reserved.

Abstract This document defines a number of names for common services, roles and functions found in SIP-

based voice-over-IP networks. It extends RFC 2142.

1 Terminology

In this document, the key words “must”, “must not”, “required”, “shall”, “shall not”, “should”, “should not”, “recommended”, “may”, and “optional” are to be interpreted as described in RFC 2119 [1] and indicate requirement levels for compliant implementations.

2 Common Functional Names

The names below should be reachable by SMTP (email), instant messaging [2] as well as SIP [3], in both lower-case, upper-case and combinations thereof. (For example, variations such as Sos, SoS, SOS should all be valid.) This extends the names enumerated in RFC 2142 [4].

sos  emergency services for members of this organization [5]
911  alias for “sos”
112  alias for “sos”
operator voice operator for organization
sipmaster contact for technical issues with SIP services
voicemail voicemail address (see also RFC 3087 [6])

TBD: Conferences?

INTERNET-DRAFT  draft-schulzrinne-sip-urls-00.ps  April 22, 2001

3 Functional Designations

sales  sales department
info  general, typically automated, information
help  general help, e.g., with locating other resources within an organization

support marketing

See Section 3, RFC 2142.

Any functional name designated by an email aliases should also be available via SIP. Local language versions of these names should also be defined.

4 Naming Users

It is recommended that organizations allow “first.last” and “first.m.last” (where “m” is the middle initial) as aliases for their users, with variations that differ only in the case of the letters, such as “First.Last” or “FIRST.M.LAST” considered equivalent to the canonical all-lower-case form.

References

[1] S. Bradner, “Key words for use in RFCs to indicate requirement levels,” Request for Comments 2119, Internet Engineering Task Force, Mar. 1997.

[2] M. Day, J. Rosenberg, and H. Sugano, “A model for presence and instant messaging,” Request for Comments 2778, Internet Engineering Task Force, Feb. 2000.

[3] M. Handley, H. Schulzrinne, E. Schooler, and J. Rosenberg, “SIP: session initiation protocol,” Request for Comments 2543, Internet Engineering Task Force, Mar. 1999.

[4] D. Crocker, “Mailbox names for common services, roles and functions,” Request for Comments 2142, Internet Engineering Task Force, May 1997.

[5] H. Schulzrinne, “Providing emergency call services for SIP-based internet telephony,” Internet Draft, Internet Engineering Task Force, Mar. 2001. Work in progress.

[6] B. Campbell and R. Sparks, “Control of service context using SIP Request-URI,” Request for Comments 3087, Internet Engineering Task Force, Apr. 2001.

5 Acknowledgements

TBD.

6 Authors’ Addresses

Henning Schulzrinne Dept. of Computer Science Columbia University 1214 Amsterdam Avenue New York, NY 10027 USA electronic mail: schulzrinne@cs.columbia.edu

Full Copyright Statement

Copyright (c) The Internet Society (2001). All Rights Reserved.

This document and translations of it may be copied and furnished to others, and derivative works that comment on or otherwise explain it or assist in its implementation may be prepared, copied, published and distributed, in whole or in part, without restriction of any kind, provided that the above copyright notice and this paragraph are included on all such copies and derivative works. However, this document itself may not be modified in any way, such as by removing the copyright notice or references to the Internet Society or other Internet organizations, except as needed for the purpose of developing Internet standards in which case the procedures for copyrights defined in the Internet Standards process must be followed, or as required to translate it into languages other than English.

The limited permissions granted above are perpetual and will not be revoked by the Internet Society or its successors or assigns.

This document and the information contained herein is provided on an ”AS IS” basis and THE IN-TERNET SOCIETY AND THE INTERNET ENGINEERING TASK FORCE DISCLAIMS ALL WAR-RANTIES, EXPRESS OR IMPLIED, INCLUDING BUT NOT LIMITED TO ANY WARRANTY THAT THE USE OF THE INFORMATION HEREIN WILL NOT INFRINGE ANY RIGHTS OR ANY IMPLIED WARRANTIES OF MERCHANTABILITY OR FITNESS FOR A PARTICULAR PURPOSE.